OpenStudy (anonymous):

find the area of the region between y=x, y=x^2+1, x=-1 x=2

OpenStudy (anonymous):

hi yeah i said the question wrong last time sorry lol.

OpenStudy (campbell_st):

ok... so same thing \[\int\limits_{-1}^{0} (x^2 + 1)dx + \frac{1}{2} \times 1 \times 1 + \int\limits_{0}^{2}(x^2 + 1)dx - \int\limits_{0}^{2} (x) dx\] that would be my solution

OpenStudy (anonymous):

hmm. okay cause we were never told to find areas of shapes in this lesson.

OpenStudy (anonymous):

thank you tho appreciate your help!

OpenStudy (campbell_st):

ok... so you for the area below the x-axis its \[\left| \int\limits_{-1}^{0} (x) dx\right|\] if you calculate the area below the axis is will be a negative... so to eliminate that problem have the integral inside absolute value symbols

OpenStudy (campbell_st):

I would have given you extra marks for using a simple method for finding an area...

OpenStudy (anonymous):

well we were taught the vertical rectangle method so i think if i use that the answer will be different than your method ? D;

OpenStudy (anonymous):

im going to see what i come up with hang on.

OpenStudy (campbell_st):

here is a geogebra worksheet with the information move the slider to get it to work. you may have to download the free software... it's worth it I have also attached the image

OpenStudy (anonymous):

I got the answer to be 9/2..

OpenStudy (anonymous):

the method you showed me gave me 11/6 -.-

OpenStudy (anonymous):

what should i do lol.

OpenStudy (anonymous):

can i show you the method i did ?

OpenStudy (anonymous):

2 ∫ x^2 + 1 - x dx = -1 ...........................2 x^3/3 + x - x^2/2 | = ..........................-1 8/3 + 2 - 2 - (-1/3 - 1 - 1/2) = 8/3 + 1/3 + 1 + 1/2 = 10 1/2

OpenStudy (anonymous):

oops thats supposed to be an integral

OpenStudy (campbell_st):

ok... the value is 4.5 so this is what I did \[\int\limits_{-1}^{0} (x^2 + 1)dx = [\frac{1}{3}x^3 + x]^{0}_{-1} = \frac{4}{3}\] and \[\left| \int\limits_{-1}^{0} (x) dx\right| = [\frac{x^2}{2}]^0_{-1} = \frac{1}{2}\] next its \[\int\limits_{0}^{2}(x^2 + 1)dx = [\frac{1}{3} x^3 + x]^{2}_{0} = \frac{8}{3}\] and lastly \[\int\limits_{0}^{2}(x) dx = [\frac{1}{2} x^2]^2_{0} = 2\] so the area between the curves is \[A = \frac{4}{3} + \frac{1}{2} + (\frac{8}{3} + 2) - 2\] I think you'ff find its \[\frac{27}{6}...or... 4\frac{1}{2}\]

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

Thank you it does!!

OpenStudy (anonymous):

Actually you know what we both got the same answer with different method, i just integrated wrong, Lmao it is 4.5 thank you!